# Connect the Dots: Knowledge Graph–Guided Crawler Attack on Retrieval-Augmented Generation Systems

### v1

@@ 5.3 Strategy Scheduler
Figure 6: The workflow of Strategy Scheduler. The scheduler selects an action from the graph $\mathcal{G}_{t}$ via a two-stage process. First, it samples an entity ($e_{t}^{*}$) based on a score that estimates CMG by balancing two key metrics. Second, it selects a relation ($r_{t}^{*}$) by identifying the entity’s largest local information deficit. A score cache with lazy updates ensures the efficiency of this process.
This module address the second challenge of the ASCP framework: attacker action spaces are intractable. Operating on the knowledge graph from the KG-Constructor, its task is to select the optimal entity-relation pair (e.g., ([Uncaptioned image], Family Medical History)) to guide the next query. The primary difficulty is that the space of potential strategic actions is combinatorially vast, rendering a brute-force search for the optimal pair computationally infeasible. Our key insight is an asymmetry between entities and relations: entities are the primary drivers of coverage expansion, while relations refine the exploration direction. Exploiting this asymmetry, we reframe the problem into a hierarchical two-stage decision: first select an entity, then pick which of its unexplored relations to probe. The workflow is shown in Fig. 6.
Entities Selection. To identify the entity that maximizes expected coverage, RAGCrawler fist estimate the CMG of each entity in $\mathcal{G}_{t}$ using two complementary metrics: empirical payoff and graph prior. The empirical payoff captures historical information gain and reflects how informative an entity has been in past queries. The graph prior exploits the topology of the knowledge graph to identify entities located in structurally under-explored regions. By jointly capturing exploitation through past utility and exploration through structural potential, these metrics provide a more complete CMG estimation. Their values are combined via a dynamically weighted sum to produce a final score for each entity, and the anchor entity $e_{t}^{*}$ is selected using robust Top-K softmax sampling.
@@
where $\bar{g}_{e}$ is the empirical information gain of $e$, $n_{e}$ is the number of times $e$ has been selected as the anchor, $N$ is the total number of anchor selections, and $c>0$ controls the strength of the confidence bonus.
The first term $\bar{g}_{e}$ estimates the contribution of $e$ to past knowledge-graph expansion. It is calculated as the normalized average increase in newly discovered entities and relations within a sliding window $t_{w}$:
$\bar{g}_{e}=\frac{1}{n_{e}}\sum_{i\in\mathcal{I}(e)}\Big(\underbrace{\Delta\mathcal{E}_{i}/\max(\Delta\mathcal{E}_{t_{w}})}_{\mathrm{normalizednewentities}}+\underbrace{\Delta\mathcal{R}_{i}/\max(\Delta\mathcal{R}_{t_{w}})}_{\mathrm{normalizednewrelations}}\Big),$ (7)
where $\mathcal{I}(e)$ denotes the set of rounds where $e$ served as the anchor, and $\Delta\mathcal{E}_{i}$ and $\Delta\mathcal{R}_{i}$ represent the newly discovered entities and relations in round $i$.
The second component is an exploration bonus, rooted in the “optimism in the face of uncertainty” principle from multi-armed bandit literature [kaufmann2012bayesian, auer2002finite]. This term adds a confidence-based reward that encourages the agent to visit less-frequently selected entities, thereby mitigating the risk of converging to a local optimum.
Graph prior. While the empirical payoff provides a self-contained mechanism for balancing exploitation and statistical exploration, it remains blind to the underlying topology of the knowledge graph. To overcome this limitation, we introduce a complementary, structure-aware exploration term: the graph prior. This prior injects topological intelligence into the strategy, directing it toward graph regions with high discovery potential. It is composed of two distinct terms:
@@
$\mathrm{AdjScore}(e)=\max_{r\in\mathcal{R}}\mathrm{Deficit}(e,r),$ (10)
@@
Here, $\mathrm{EdgeType}(e)$ is the set of relation types already connected to $e$, and $\mathcal{E}_{e.\mathrm{type}}$ is the set of all entities sharing the same semantic type as $e$.
Entity scoring and sampling. We integrate empirical payoff and graph prior into a final composite score. Critically, their relative importance may change over the course of an attack. We therefore combine them using time-varying weights to enable dynamic control of the exploration-exploitation balance:
$\mathrm{Score}(e)\;=\;\alpha_{t}\cdot\mathrm{EmpricalPayoff}(e)\;+\;\beta_{t}\cdot\mathrm{GraphPrior}(e),$ (12)
@@
Finally, to mitigate the risk of prematurely converging on a single, seemingly optimal path due to noisy gain estimates, we avoid a deterministic argmax selection. Instead, we sample the anchor entity $e_{t}^{*}$ from a Top-$K$ softmax distribution over the candidate scores, promoting strategic diversity and enhancing robustness.
@@
$r^{*}(e_{t}^{*})\;=\;\arg\max_{r\in\mathcal{R}\setminus\mathrm{EdgeType}(e_{t}^{*})}\ \mathrm{Deficit}(e_{t}^{*},r).$ (13)
@@
$\mathcal{DS}_{t}\;=\;\big\{\,\mathrm{Deficit}(e,r)\;:\;e\in\mathcal{E}_{t},\ r\in\mathcal{R}\setminus\mathrm{EdgeType}(e)\,\big\},$ (14)
@@
Optimization. To ensure scalability, the scheduler employs a lightweight caching mechanism. Previously computed $\mathrm{Score}(e)$ values are stored in a score cache. Following an update to the knowledge graph $\mathcal{G}_{t}$, a lazy-update strategy is triggered: only the scores of entities directly affected by the changes are invalidated and recomputed. The scores for all unaffected entities are served directly from the cache. This approach decouples the computational cost of the scheduling step from the total size of the graph, ensuring that the decision loop remains efficient even as the attacker’s graph expands significantly.
@@ 5.4 Query Generator
Figure 7: The workflow of Query Generator. It receives an anchor pair $(e_{t}^{*},r_{t}^{*})$ and selects a generation strategy. A candidate query $q_{t+1}$ is produced and checked against historical queries. If too similar, the generator requests a new anchor from the scheduler and resamples.
This module addresses the third challenge in the ASCP formulation: Real-world RAG systems impose practical constraints. We cannot directly send a structured anchor $(e_{t}^{*},r_{t}^{*})$ to the system, and relying solely on simple templates is not a viable solution, as it can easily be flagged by query provenance mechanisms. Therefore, the core task of this module is to translate the abstract, strategic anchor pair from the Strategy Scheduler into a concrete, executable, and natural-sounding query $q_{t+1}$. This process must address two primary challenges: 1) generating a query that is contextually relevant to the anchor’s strategic intent, and 2) ensuring the query remains novel to avoid redundancy and maximize the utility of the limited query budget.
To achieve this, the generator involves three key steps: 1) adaptive query formulation, 2) history-aware de-duplication, and 3) a feedback mechanism for system-level learning. The workflow is shown in Fig. 7.
@@
2) History-aware De-duplication To maximize the information yield of each query, we introduce a robust quality control gate powered by a history-aware resampling loop. Before being dispatched, each candidate query $q$ is compared against the historical query log $\mathcal{Q}_{\mathrm{hist}}$. If its similarity $\mathrm{sim}(q,\mathcal{Q}_{\mathrm{hist}})\!\geq\!\tau$, the generator triggers a resample operation, drawing a new anchor from the Strategy Scheduler’s Top-$K$ distribution and re-attempting the formulation, up to $M$ trials. This mechanism is critical for avoiding diminishing returns. Furthermore, this loop provides an elegant convergence heuristic: if all $M$ trials yield duplicate queries, we declare that the system has reached global convergence.
3) Penalties and Feedback Loop. Beyond producing a valid query, the generator’s final role is to enable system-level learning. Critically, queries result in refusals, empty responses or identified as duplicates incur strategy-specific penalties. This penalty signal is fed back to the scheduler as a reduced payoff, directly influencing its subsequent Empirical Payoff score calculations and anchor selections. This feedback mechanism, which connects the outcome of the generator’s action back to the high-level strategy, allows the system to learn from failed interactions and dynamically adapt its strategy over time.
@@ 6 Evaluation
  • RQ1: How effective is RAGCrawler compared with existing methods?
  • RQ2: How does the the retriever in victim RAG affect coverage performance?
  • RQ3: How does the internal LLM agent within attack method influence performance?
@@
  • RQ5: How do the hyperparameters and each technique affect effectiveness of RAGCrawler?
@@ 6.1 Evaluation Setup
Dataset. We evaluate on four datasets spanning scientific and medical domains with varying styles and scales. Three datasets, TREC-COVID, SciDocs, and NFCorpus, are drawn from the BEIR benchmark [thakur2021beir], containing about 171.3K, 25.6K, and 5.4K documents, respectively. We further include Healthcare-Magic-100k [healthcaremagic] (abbreviated as Healthcare), with roughly 100K patient–provider Q&A samples. TREC-COVID, SciDocs, NFCorpus, and Healthcare focus on biomedical literature, scientific papers, consumer health, and clinical dialogues, respectively. For efficiency and fair comparison, we randomly select 1,000 de-duplicated documents from each corpus as the victim RAG collection [jiang2024rag, di2024pirates]. We confirm that the sampled subsets preserve similar semantic distributions to the full corpora using Energy Distance [szekely2013energy] and the Classifier Two-Sample Test (C2ST) [lopez2016revisiting], with details in Appendix 9.1.
@@
Generator and Retriever. We employ two retrievers in our evaluations: BGE [zhang2023retrieve] and GTE [li2023towards]. For generators, we evaluate four large language models: Llama 3.1 Instruct-8B (denoted as Llama-3-8B) [dubey2024llama], Command-R-7B [command-r7b], Microsoft-Phi-4 [abdin2024phi], and GPT-4o-mini [gpt4o-mini]. These generators represent diverse families of reasoning and instruction-tuned models, covering both open-source and proprietary systems and varying parameter scales. The inclusion of GPT-4o-mini, with its built-in safety guardrails [gpt-guardrail], demonstrates our attack’s robustness against industry defenses.
@@
Attacker-side LLM. We adopt two attacker-side LLMs to simulate adversarial query generation. In the main experiments, we use Doubao-1.5-lite-32K [doubao] due to its high speed, cost efficiency, and independence from the RAG’s generator family. This aligns with the black-box assumption. We also evaluate a smaller open-source alternative, Qwen-2.5-7B-Instruct [qwen2.5-7b], to examine transferability between different model families and sizes.
@@
RAG Setting. We cover three RAG configurations: vanilla RAG, RAG with query rewriting, and RAG employing multi-query retrieval with RRF re-ranking (default: 3 queries per request). We set the retrieval number to $k=10$ and analyze its impact in Sec. 6.6. The retrieved documents are provided to the generator as contextual input through a structured system prompt. Further details on the system prompts are provided in Appendix 9.5.
@@
Baselines. We compare RAGCrawler against two state-of-the-art black-box RAG extraction attacks: RAGThief [jiang2024rag] and IKEA [wang2025silent]. RAGThief represents continuation-based attack, and IKEA represents keyword-based attack. To ensure a fair comparison, all attacks employ the same attacker-side LLM. For all experiments, we use the default hyperparameters reported in the original papers. Each attack is allowed to issue at most 1,000 queries to the victim RAG, ensuring consistent query budgets across methods.
Metrics. We evaluate extraction along two axes coverage and content quality: (i) Coverage Rate. This metric quantifies how much of the victim RAG’s private corpus has been exposed through retrieval, corresponding to the adversary’s goal in Eq. 3. Following prior work [jiang2024rag, wang2025silent], this metric is only calculated on queries for which the victim RAG produces a non-refusal response. (ii) Semantic Fidelity. To quantify the semantic overlap between the private corpus $\mathcal{D}$ and the extracted snippets $\mathcal{A}$, we first measure the similarity for each document $d\in\mathcal{D}$ to its best-aligned snippet using a fixed encoder $E_{\mathrm{eval}}(\cdot)$:
$s(d)=\max_{e\in\mathcal{E}}\ \cos\bigl(E_{\mathrm{eval}}(d),\,E_{\mathrm{eval}}(e)\bigr).$ (15)
@@
The average of $s(d)$ over the entire corpus represents the overall semantic fidelity of the extracted text. To assess practical utility, we also measure Reconstruction Fidelity. We construct a surrogate RAG from the extracted knowledge and evaluate it using the official query sets corresponding to the documents in TREC-COVID, SciDocs, and NFCorpus corpora, totaling 1,860, 2,181, and 1,991 queries, respectively. The Healthcare dataset is excluded due to the absence of official query annotations. We report: (a)Success Rate (fraction of non-refusal responses), (b) Answer Similarity (embedding cosine similarity between surrogate and victim answers); and (c) ROUGE-L score.
@@ 6.2 Effectiveness (RQ1)
We evaluate the effectiveness of RAGCrawler through three key metrics: corpus coverage rate, semantic fidelity and reconstruction fidelity. Across all evaluations, RAGCrawler demonstrates a significant leap in performance over existing baseline attacks.
TABLE I: Coverage Rate (CR) and Semantic Fidelity (SF) of attacks across datasets and generators (BGE Retriever). Best results are in bold. RAGCrawler consistently achieves the highest CR and SF, outperforming baselines across all settings.
@@
Dataset Generator Metric RAGTheif IKEA RAGCrawler TREC-COVID Llama-3-8B CR 0.131 0.161 0.494 SF 0.447 0.495 0.591 Command-R CR 0.154 0.173 0.544 SF 0.444 0.525 0.614 Microsoft-Phi-4 CR 0.121 0.197 0.474 SF 0.468 0.547 0.619 GPT-4o-mini CR 0.000 0.197 0.465 SF - 0.543 0.572 Scidocs Llama-3-8B CR 0.053 0.513 0.661 SF 0.264 0.495 0.523 Command-R CR 0.093 0.522 0.717 SF 0.295 0.514 0.561 Microsoft-Phi-4 CR 0.065 0.545 0.711 SF 0.324 0.534 0.563 GPT-4o-mini CR 0.000 0.421 0.516 SF - 0.475 0.492 NFCorpus Llama-3-8B CR 0.061 0.503 0.797 SF 0.451 0.644 0.698 Command-R CR 0.169 0.517 0.844 SF 0.467 0.656 0.705 Microsoft-Phi-4 CR 0.113 0.566 0.813 SF 0.487 0.663 0.717 GPT-4o-mini CR 0.000 0.493 0.631 SF - 0.631 0.653 HealthCare Llama-3-8B CR 0.361 0.687 0.807 SF 0.536 0.588 0.618 Command-R CR 0.170 0.592 0.766 SF 0.383 0.578 0.582 Microsoft-Phi-4 CR 0.052 0.588 0.654 SF 0.434 0.558 0.599 GPT-4o-mini CR 0.000 0.693 0.799 SF - 0.490 0.577
Figure 8: Coverage Rate vs. Query Number (1,000 Budget) across datasets and generators (BGE Retriever). RAGCrawler steadily increases coverage and consistently surpasses both RAGThief and IKEA. Comprehensive results are in Appendix 9.4.
Coverage Rate. RAGCrawler achieves consistently superior extraction efficiency and breadth. As shown in the coverage rate curves in Fig. 8 and Fig. 12, RAGCrawler steadily increases the corpus coverage throughout the attack, reaching a final average rate of 66.8% under a 1,000-query budget (Tab. I). This high efficiency is particularly pronounced in challenging, broad-domain corpora such as SciDocs and TREC-COVID. The success of RAGCrawler is attributable to its global knowledge graph and UCB-based planning algorithm. This design enables it to dynamically and strategically traverse the entire semantic space of the corpus, ensuring it continuously targets and extracts from underexplored regions. Furthermore, RAGCrawler remains highly effective against RAG systems with built-in guardrails (e.g., those using GPT-4o-mini). Its use of stealthy, policy-compliant prompts allows it to circumvent defenses that easily detect and block the explicit jailbreak attempts used by other methods.
In contrast, baseline methods exhibit significant limitations. The strongest baseline, IKEA, plateaus early and achieves an average coverage of only 46.1% (a 20.7 percentage point deficit compared to RAGCrawler). RAGThief is even less effective, with an average coverage of a mere 9.6%. These heuristic-based methods tend to repeatedly query semantically similar regions or drift off-topic, failing to achieve comprehensive corpus exploration. RAGThief, in particular, fails completely against guarded models GPT-4o-mini, as its direct jailbreak prompts are consistently refused.
Semantic Fidelity. Beyond extracting a larger volume of content, RAGCrawler also ensures the extracted information is of higher quality. RAGCrawler achieves the highest Semantic Fidelity (SF) score across all dataset-generator pairs, with a cross-dataset average of 0.605 (Tab. I). This indicates that the content it extracts is highly faithful to the source documents. This high fidelity is a direct result of our design of the Query Generator, which maintains contextual coherence and ensures that the extracted text chunks are semantically sound and accurate. By comparison, IKEA and RAGThief score lower, with average SF scores of 0.559 and 0.417, respectively. Their less-targeted query strategies may yield fragmented or contextually distorted content, diminishing the quality of the extracted knowledge.
TABLE II: Reconstruction Fidelity of Surrogate RAG Systems using extracted knowledge (Llama-3-8B Generator, BGE Retriever). Best results are in bold. Surrogates built from extracted knowledge of RAGCrawler achieve the highest performance in all metrics.

### v2

@@ 6.3 Strategy Scheduler
This module addresses the second challenge of the ASCP formulation: attacker action space is intractable. Operating on the knowledge graph from the KG-Constructor, its task is to select the optimal entity-relation pair (e.g., ( , HotFix)) to guide the next query. The challenge is the space of potential strategic actions is combinatorially vast, rendering a brute-force search for the optimal pair computationally infeasible.
Our key insight is an asymmetry between entities and relations: entities are the primary drivers of coverage expansion, while relations refine the exploration direction. Exploiting this asymmetry, we reframe the problem into a hierarchical two-stage decision: first select an entity, then pick its unexplored relation to probe. The workflow is shown in Fig. 6.
Figure 6: Strategy Scheduler. The scheduler selects an action from the graph $\mathcal{G}_{t}$ via a two-stage process. First, it samples an entity ($e_{t}^{*}$) based on a score that estimates CMG by balancing two key metrics. Second, it selects a relation ($r_{t}^{*}$) by identifying the entity’s largest local information deficit. A score cache with lazy updates ensures the efficiency of this process.
Entities Selection. We estimate the utility of each entity in $\mathcal{G}_{t}$ using two complementary metrics: empirical payoff and graph prior. The empirical payoff captures historical information gain and reflects how informative an entity has been in past queries. The graph prior exploits the topology of the knowledge graph to identify entities located in structurally under-explored regions. We combine them into a composite score and sample the anchor entity from a Top-$K$ softmax distribution to preserve diversity under noisy estimates.
@@
where $n_{e}$ is the number of times $e$ has been selected as the anchor, $N$ is the total number of anchor selections so far, and $c>0$ controls the strength of the bonus [kaufmann2012bayesian, auer2002finite].
We define $\bar{g}_{e}$ as the average normalized graph growth observed when selecting $e$:
$\bar{g}_{e}\;=\;\frac{1}{\max(1,n_{e})}\sum_{i\in\mathcal{I}(e)}\left(\frac{\Delta\mathcal{E}_{i}}{\Delta\mathcal{E}_{\max}+\epsilon}+\frac{\Delta\mathcal{R}_{i}}{\Delta\mathcal{R}_{\max}+\epsilon}\right),$ (9)
where $\mathcal{I}(e)$ is the set of rounds in which $e$ served as the anchor, and $\Delta\mathcal{E}_{i}$ and $\Delta\mathcal{R}_{i}$ are the numbers of newly added entities and relation types in round $i$ after semantic merging. To make gains comparable across rounds, we normalize by window maxima computed over the most recent $t_{w}$ rounds, denoted by $\mathcal{W}_{t}$, i.e., $\Delta\mathcal{E}_{\max}=\max_{j\in\mathcal{W}_{t}}\Delta\mathcal{E}_{j}$ and $\Delta\mathcal{R}_{\max}=\max_{j\in\mathcal{W}_{t}}\Delta\mathcal{R}_{j}$.
Graph prior. While the empirical payoff provides a self-contained mechanism for balancing exploitation and statistical exploration, it remains blind to the underlying topology of the knowledge graph. Therefore, we introduce a complementary, structure-aware exploration term: the graph prior. This prior injects topological intelligence into the strategy, directing it toward graph regions with high discovery potential. It is composed of two distinct terms:
@@
$\mathrm{AdjScore}(e)=\max_{r\in\mathcal{R}_{t}}\mathrm{Deficit}(e,r),$ (12)
@@
Here, $\mathcal{R}_{t}$ denotes the set of relation types currently in the graph $\mathcal{G}_{t}$, $\mathrm{EdgeType}(e)$ is the set of relation types already connected to $e$, and $\mathcal{E}_{e.\mathrm{type}}$ is the set of all entities sharing the same semantic type as $e$.
Entity scoring and sampling. We integrate empirical payoff and graph prior into a final composite score using time-varying weights:
$\mathrm{Score}(e)\;=\;\alpha_{t}\cdot\mathrm{EmpiricalPayoff}(e)\;+\;\beta_{t}\cdot\mathrm{GraphPrior}(e),$ (14)
@@
To mitigate the risk of prematurely converging on a seemingly optimal path due to noisy gain estimates, we sample the anchor entity $e_{t}^{*}$ from a Top-$K$ softmax distribution over the candidate scores, enhancing strategic diversity and robustness.
@@
$r^{*}(e_{t}^{*})\;=\;\arg\max_{r\in\mathcal{R}_{t}\setminus\mathrm{EdgeType}(e_{t}^{*})}\ \mathrm{Deficit}(e_{t}^{*},r).$ (15)
@@
$\mathcal{DS}_{t}\;=\;\big\{\,\mathrm{Deficit}(e,r)\;:\;e\in\mathcal{E}_{t},\ r\in\mathcal{R}_{t}\setminus\mathrm{EdgeType}(e)\,\big\},$ (16)
@@
Optimization. We keep scheduling scalable by caching entity scores and lazily recomputing only affected entries after each graph update. When $\mathcal{G}_{t}$ is updated, we invalidate the cache for entities whose neighborhoods changed, and we reuse cached scores for all other entities. This reduces per-round overhead and decouples scheduling cost from the full graph size.
@@ 6.4 Query Generator
This module addresses the third challenge in the ASCP formulation: Real-world RAG systems impose practical constraints. We cannot directly send a structured anchor $Anc_{t}$ to the system, and relying solely on simple templates is not a viable solution, as it can easily be flagged by query provenance mechanisms. Therefore, the core task of this module is to translate the abstract, strategic anchor pair from the Strategy Scheduler into a concrete, executable, and natural-sounding query $q_{t+1}$.
This process must address two primary challenges: 1) generating a query that is contextually relevant to the anchor’s strategic intent, and 2) ensuring the query remains novel to avoid redundancy under the limited query budget. To achieve this, the generator (Fig. 7) involves the following three steps.
Figure 7: Query Generator. Given an anchor pair $(e_{t}^{*},r_{t}^{*})$, it selects a generation strategy. A candidate query $q_{t+1}$ is produced and checked against historical queries. If redundant, it requests a new anchor from the scheduler and resamples.
@@
2) History-aware De-duplication To maximize the information yield of each query, we introduce a robust quality control gate powered by a history-aware resampling loop. Before being dispatched, each candidate query $q$ is compared against the historical query log $\mathcal{Q}_{\mathrm{hist}}$. If its similarity $\mathrm{sim}(q,\mathcal{Q}_{\mathrm{hist}})\!\geq\!\tau_{dup}$, the generator triggers a resample operation, drawing a new anchor from the Strategy Scheduler’s Top-$K$ distribution and re-attempting the formulation, up to $M$ trials. This mechanism is critical for avoiding diminishing returns. If all $M$ trials yield near-duplicate queries, we treat the current exploration as saturated. As a practical fallback, an attacker can either restart a seeding step to diversify the query pool [wang2025silent, jiang2024rag], or terminate early to avoid wasting the remaining query budget.
3) Penalties and Feedback Loop. We close the loop by feeding generation outcomes back into scheduling. When a query is refused, yields an empty answer, or is rejected as a duplicate, we apply a strategy-specific penalty to the corresponding payoff update so that such anchors become less likely to be selected in future rounds. This feedback lets the planner adapt away from unproductive regions and toward anchors that reliably produce new graph growth.
@@ 7 Evaluation
RQ1: How effective is RAGCrawler compared with existing attack methods?
RQ2: How does the retriever in the victim RAG affect stealing performance?
RQ3: How does the attacker-side LLM agent influence stealing performance?
@@
RQ5: How do hyperparameters and individual techniques affect the effectiveness of RAGCrawler?
@@ 7.1 Evaluation Setup
Dataset. We evaluate on four corpora spanning diverse domains with varying styles and scales. TREC-COVID and SciDocs are from BEIR [thakur2021beir] with 171.3K and 25.6K documents, and NFCorpus [boteva2016full] with 5.4K documents. We additionally include Healthcare-Magic-100k [healthcaremagic] (Healthcare) with about 100K patient-provider Q&A samples. Together they cover biomedical literature, scientific papers, consumer health, and clinical dialogues. For efficiency and fair comparison, we randomly sample 1,000 de-duplicated documents from each corpus to form the victim RAG collection [jiang2024rag, di2024pirates]. We verify that each subset preserves the full-corpus semantic distribution using Energy Distance [szekely2013energy] and C2ST [lopez2016revisiting] (Appendix B), and report results with larger victim collections in Appendix E.4.
@@
Generator and Retriever. We employ two retrievers in our evaluations: BGE [zhang2023retrieve] and GTE [li2023towards]. For generators, we evaluate four large language models: Llama 3.1 Instruct-8B (denoted as Llama-3-8B) [dubey2024llama], Command-R-7B [command-r7b], Microsoft-Phi-4 [abdin2024phi], and GPT-4o-mini [gpt4o-mini]. These generators cover both open-source and proprietary systems and span different model families and sizes. Including GPT-4o-mini, which incorporates safety guardrails [gpt-guardrail], allows us to evaluate whether stealing remains feasible under built-in guardrails.
@@
Attacker-side LLM. We adopt two attacker-side LLMs to simulate query generation. In the main experiments, we use Doubao-1.5-lite-32K [doubao] due to its high speed, cost efficiency, and independence from the RAG’s generator family (aligns with the black-box assumption). We also evaluate a smaller open-source alternative, Qwen-2.5-7B-Instruct [qwen2.5-7b], to examine transferability between model families and sizes.
@@
RAG Setting. We consider three victim configurations: vanilla RAG, query rewriting, and multi-query retrieval with RRF re-ranking (default: 3 query variants per user request). We set retrieval depth to $k=10$ by default and analyze the impact of $k$ in Sec. 7.6. The retrieved documents are provided to the generator as contextual input through a structured system prompt. Further details are provided in Appendix F.
@@
Baselines. We compare RAGCrawler against two state-of-the-art black-box RAG knowledge-base stealing attacks: RAGThief [jiang2024rag] and IKEA [wang2025silent]. RAGThief represents continuation-based attack, and IKEA represents keyword-based attack. To ensure fair comparison, all attacks use the same attacker-side LLM, and we adopt default hyperparameters from the original papers. Unless otherwise specified, each attack is allowed to issue at most 1,000 queries to the victim RAG, ensuring consistent query budgets across methods.
Figure 8: Coverage Rate vs. Query Number (1,000 Budget) across datasets and generators (BGE Retriever). RAGCrawler steadily increases coverage and consistently surpasses both RAGThief and IKEA. Full results are in Fig. 11.
Metrics. We evaluate stealing along four complementary axes: exposure breadth, semantic faithfulness of the stolen content, query cost to reach target coverage, and whether the extracted corpus can support service replication.
Coverage Rate (CR) measures the fraction of the corpus exposed through retrieval, matching Eq. 3; refusals are treated as empty outcomes as in [jiang2024rag, wang2025silent].
Semantic Fidelity (SF) measures average semantic overlap between each document $d\in\mathcal{D}$ and the stolen content $\mathcal{A}$, where $\mathcal{A}$ is the set of answer snippets produced across all issued queries, using a fixed encoder $E_{\mathrm{eval}}(\cdot)$:
$s(d)=\max_{e\in\mathcal{A}}\ \cos\bigl(E_{\mathrm{eval}}(d),\,E_{\mathrm{eval}}(e)\bigr).$ (17)
Target-Coverage Query Complexity $Q_{\gamma}$ is the minimum number of issued queries needed to reach coverage $\gamma$ (the first $t$ such that $\mathrm{CR}(t)\geq\gamma$). When the target is not reached under a cap $Q_{\max}$, we treat $Q_{\gamma}$ as right-censored and report $Q_{\gamma}>Q_{\max}$.
@@
Reconstruction Fidelity assesses whether the stolen corpus supports service replication. We build a surrogate RAG from the stolen content and evaluate it on official query annotations of TREC-COVID, SciDocs, and NFCorpus, restricted to the sampled documents (1,860, 2,181, and 1,991 queries). Healthcare is excluded due to the lack of official query annotations. We report success rate (fraction of non-refusal responses), answer similarity (semantic similarity between surrogate and victim answers), and ROUGE-L.
@@ 7.2 Effectiveness (RQ1)
We evaluate the effectiveness of RAGCrawler through corpus coverage, semantic fidelity, target-coverage query complexity, and reconstruction fidelity. Across all evaluations, RAGCrawler demonstrates the strongest performance.
Table 1: Coverage rate (CR) and semantic fidelity (SF) of attacks under a 1,000-query budget (victim retriever: BGE) across four datasets and four generators. Abbreviations: Gen (Generator), TREC-C (TREC-COVID), Cmd-R (Command-R), MS-Phi (Microsoft-Phi-4), 4o-mini (GPT-4o-mini). “N/A” indicates undefined semantic fidelity due to zero coverage. RAGCrawler performs best in every setting.
@@
Dataset Gen CR SF RAGThief IKEA RAGCrawler RAGThief IKEA RAGCrawler TREC-C Llama 0.131 0.161 0.494 0.447 0.495 0.591 Cmd-R 0.154 0.173 0.544 0.444 0.525 0.614 MS-Phi 0.121 0.197 0.474 0.468 0.547 0.619 4o-mini 0.000 0.197 0.465 N/A 0.543 0.572 SciDocs Llama 0.053 0.513 0.661 0.264 0.495 0.523 Cmd-R 0.093 0.522 0.717 0.295 0.514 0.561 MS-Phi 0.065 0.545 0.711 0.324 0.534 0.563 4o-mini 0.000 0.421 0.516 N/A 0.475 0.492 NFCorpus Llama 0.061 0.503 0.797 0.451 0.644 0.698 Cmd-R 0.169 0.517 0.844 0.467 0.656 0.705 MS-Phi 0.113 0.566 0.813 0.487 0.663 0.717 4o-mini 0.000 0.493 0.631 N/A 0.631 0.653 Healthcare Llama 0.361 0.687 0.807 0.536 0.588 0.618 Cmd-R 0.170 0.592 0.766 0.383 0.578 0.582 MS-Phi 0.052 0.588 0.654 0.434 0.558 0.599 4o-mini 0.000 0.693 0.799 N/A 0.490 0.577 Average 0.096 0.461 0.668 0.417 0.559 0.605
Coverage Rate. Under a fixed 1,000-query budget with the BGE retriever, RAGCrawler attains the highest final coverage in every configuration in Table 1. Averaged over all 16 configurations, RAGCrawler exposes 66.8% of the corpus, compared with 46.1% for IKEA and 9.6% for RAGThief. The margin is largest on harder corpora (TREC-COVID), where RAGCrawler reaches 49.4% average coverage across generators, while IKEA remains at 18.2%. The coverage curves in Fig. 8 further indicate that RAGCrawler continues to accrue new documents throughout the process, whereas baselines typically flatten earlier, consistent with diminishing returns from locally reactive querying.
RAGCrawler also remains effective under generators with built-in guardrails. When the victim generator is GPT-4o-mini, RAGThief yields zero coverage across all four corpora, while RAGCrawler still reaches 46.5%-79.9% coverage, suggesting that benign, policy-compliant probing can sustain stealing even when overt jailbreak patterns are blocked.
Semantic Fidelity. Higher coverage does not come at the expense of semantic quality. Across all dataset–generator pairs, RAGCrawler achieves the highest average SF (0.605), exceeding IKEA (0.559) and RAGThief (0.417) in Table 1.
Table 2: Target-coverage query complexity to reach 70% coverage ($Q_{0.7}$) under a cap $Q_{\max}{=}5{,}000$ (victim: Llama-3-8B generator, BGE retriever). Entries marked $Q_{\max}$” are right-censored, with the achieved coverage shown in parentheses. The full sweep of $Q_{\gamma}$ is reported in Tab. 10. RAGCrawler reaches high coverage with far fewer queries.